Solution of : xp^2 - (x-a)^2 = 0
1
Expert's answer
2020-12-21T17:30:00-0500

"xp^2-(x-a)^2=0"

we have difference of squares

"(\sqrt{x}p-(x-a))(\sqrt{x}p+(x-a))=0"

Two equations and two solutions:

first equation:

"\sqrt{x}p-x+a=0"

"p=\frac{x-a}{\sqrt{x}}=\sqrt{x}-\frac{a}{\sqrt{x}}"

first solution:

"z=\frac{2}{3}x^{\frac{3}{2}}-2a\sqrt{x}+C"

second equation:

"\sqrt{x}p+x-a=0"

"p=-\frac{x-a}{\sqrt{x}}=-\sqrt{x}+\frac{a}{\sqrt{x}}"

second solution:

"z=-\frac{2}{3}x^{\frac{3}{2}}+2a\sqrt{x}+C"